# Sequence-aware Pre-training for Echocardiography Probe Movement Guidance

### v2

@@ 4.1 Implementation Details
Dataset. The echocardiographic scanning dataset contains 356 demonstrations from routine clinical scans, with a total of 1.31 million image-pose pairs. The dataset was collected using a GE machine (General Electric, USA) equipped with a M5S probe. In the experiment, we divided the dataset into 284 scans (1.03 million sample pairs) for the training set and 72 scans (0.28 million sample pairs) for the validation set. Please note that the validation set comprises individuals not encountered during training, allowing for an assessment of the algorithm’s generalization capabilities. The whole data collecting process was approved and supervised by The Tsinghua University Medical Ethics Committee.
@@
Pre-training. The vision encoder is firstly pre-trained on training set using the I-JEPA method [13] with their default hyper-parameters. During our pre-training, the vision encoder is further jointly trained with the sequence-aware transformer. The training batch size is 512. The model is optimized using AdamW for 50 epochs, with the first 7 epochs designated for warm-up, followed by a cosine learning rate schedule starting at an initial learning rate of 0.00025. Data augmentation only involves random resized crop, with a crop size of 224$\times$224. The mask ratio is set to range between 0.3 and 0.5. The default sequence length $L$ is set to 6 unless otherwise specified.
Downstream fine-tuning. For fine-tuning, after loading the pre-trained parameters, the model is optimized using AdamW with a batch size of 512 and a learning rate of $1\times 10^{-4}$. A cosine learning rate schedule is employed, and the training is conducted for 5 epochs. For evaluation protocols (1)/(2), the vision encoder is jointly fine-tuned with other modules. Additionally, for evaluation protocol (2), we apply a drop path rate of 0.2 to the sequence-aware transformer and a layer-wise learning rate decay of 0.4 to the vision encoder. The baselines use the same settings. For evaluation protocol (3), we empirically find that freezing the vision encoder achieves better performance.
@@
Translation MAE $=\frac{1}{3}\sum_{i=1}^{3}|\mathbf{a}_{i}-\mathbf{a}^{{}^{\prime}}_{i}|,\ \ \text{Rotation MAE}=\frac{1}{3}\sum_{i=4}^{6}|\mathbf{a}_{i}-\mathbf{a}^{{}^{\prime}}_{i}|.$ (11)
@@
[…]
@@
[…]
@@
[…]
@@ 4.3 Ablation and Discussion of Key Components
Table 2: Ablation study on sequence-aware pre-training.
Pre-train     Protocol Trans.(mm) $\downarrow$ Rot.(°) $\downarrow$     Protocol Trans.(mm) $\downarrow$ Rot.(°) $\downarrow$
✗     Unidirectional 7.41 7.90     Bidirectional 5.53 6.69
✓     7.01 (-5.4%) 7.55 (-4.4%)     5.12 (-7.4%) 6.34 (-5.2%)
@@
Protocol Trans.(mm) $\downarrow$ Rot.(°) $\downarrow$
@@
Figure 7: Ablation study under various training scans and sequence length.
Figure 8: Ablation of learning personalized features. Left: Pre-training improves personalized feature learning. Right: Sequential modeling aids in learning personalized features.
@@
Furthermore, since the heart undergoes continuous beating, which affects the content of cardiac ultrasound images, we further investigate the impact of the cardiac cycle on the model. Specifically, we sample 500 pairs of images from the same view but at different cardiac cycle phases, generate guidance actions toward ten standard views, and compute the average absolute error/difference between outputs from different cardiac phases. For the sequence modeling approach, we keep the same $L-1$ historical views unchanged and only varied the current input’s cardiac phase. As shown in Fig. 9, the sequential modeling method significantly improves robustness against cardiac cycle variations, whereas the single-image-based approach is heavily influenced by the cardiac phase.
@@
Table 4: Ablation study on sampling strategy. MAE values are averages of ten planes.
Strategy     Protocol Trans.(mm) $\downarrow$ Rot.(°) $\downarrow$     Protocol Trans.(mm) $\downarrow$ Rot.(°) $\downarrow$
@@
Sampling strategy. Since the scanning process generates long historical trajectories, selecting key instances from these trajectories is crucial for the model’s performance. In Tab. 4, we compare segmental sampling with two common sampling strategies: dense sampling [34] and random sampling. Dense sampling selects a sequence of $L-1$ consecutive frames preceding the current image, while random sampling selects $L-1$ frames randomly from the sampling interval to form a sequence. From the table, it can be observed that under both protocols, segmental sampling achieves the best performance, while dense sampling performs the worst, and random sampling is slightly better than dense sampling. Dense sampling performs poorly because the sampled interval is too localized, and the variation in probe poses within such a local interval is typically small. This means that the sequence contains highly localized cardiac structural information around the current plane, failing to provide global structural insights which is useful for navigating to other planes. Meanwhile, the sequences obtained through random sampling can cover a larger sampling interval, resulting in a greater variety of probe poses (spanning multiple different planes), which also means more diverse structural information. However, due to its randomness, random sampling may also face the same issue as dense sampling, where the samples could be too localized. Unlike the previous two methods, segmental sampling first divides the sampling interval into $L-1$ equal segments and then randomly samples one instance from each segment. This approach prevents samples from being confined to localized intervals, which in turn leads to an improvement in overall performance.
@@ 4.4 Ablation of Mask Hyper-parameters
Figure 11: Ablations of mask type and ratio. MAE values are averages of ten planes.
@@
Figure 12: Visualization of the model’s prediction. We performed nearest-neighbor sampling from the scans to retrieve the plane most close to the model’s predictions.
@@ 4.5 Visualization
To intuitively demonstrate the model’s output, we calculate the final plane pose based on the current plane’s pose and predicted probe movement action. Using this pose, we perform nearest-neighbor retrieval from the same scan and present retrieved planes. We compare four models in Fig. 12, showing that our model’s preditions align closest with the target plane, confirming the superior precision of our guidance signal.
@@ 4.6 Efficiency
In ultrasound scanning, highly skilled sonographers make immediate adjustments based on the ultrasound images, a process that involves real-time decision-making. Therefore, an ultrasound guidance model must also possess real-time decision-making capabilities to provide timely guidance. To this end, we benchmarked inference latency on both RTX 3090 and A100 GPUs. With sequence length=6, processing times were 10.6ms (RTX 3090) and 4.9ms (A100), enabling real-time guidance.
@@ 5 Discussion and Conclusion
In this paper, we propose a cardiac probe movement guidance model which has the potential of driving autonomous robotic echocardiography or helping novice in scanning ten important standard planes. Specifically, we introduce a novel sequence-aware pre-training method to enhance the model’s ability to capture personalized cardiac structures from past sequences, addressing individual variability. Extensive experiments on a large-scale expert scanning dataset with 1.31M samples demonstrate that our model outperforms other advanced baselines and highlights the superiority of the proposed sequence-aware pre-training method.
Although experiments have validated the effectiveness of our idea, it should be noted that there still exist two limitations. Firstly, the sampling strategy is manually designed, which may be a suboptimal approach. As evidenced by the ablation study, as the sequence length increases, the rate of error reduction slows down, which may be due to the presence of information redundancy or even noise in the sampled instances. Therefore, a valuable future direction is to develop methods for sampling the most informative sequences, maximizing information gain from extensive past scanning sequence data. Secondly, information from other modalities such as ultrasound reports and sonographers’ gaze information has not yet been collected. However, this type of data has the potential to further enhance the model’s performance. We believe these limitations are also promising research directions that are worth further investigation.

### v3

@@ 4.1 Implementation Details
Dataset. The echocardiographic scanning dataset contains 476 demonstrations from routine clinical scans, with a total of 1.67 million image-pose pairs. The dataset was collected using a GE Vivid E7 machine (General Electric, USA) equipped with a M5S probe. In the experiments, we divided the dataset into 284 scans (1.03 million sample pairs) for the training set, 72 scans (0.28 million sample pairs) for validation set 1, and 120 scans for validation set 2 (0.36 million sample pairs). Validation set 1 includes 38 subjects with an age distribution of 41.1 ± 15.3 years (max: 69, min: 18). Among them, 9 subjects have cardiac diseases, including atrial septal defect, ventricular septal defect, degenerative valve changes, and ventricular septal hypertrophy. The oldest patient in this subgroup is 69 years old, with a mean age of 47.5 years. Validation set 2 consists of 60 subjects with an age distribution of 54.2 ± 10.7 years (max: 75, min: 26). The cohort was divided into three age groups: under 50 years, 50 to 60 years, and over 60 years, with each group containing 20 subjects and maintaining a 1:1 male-to-female ratio. Unless otherwise specified, experiments were conducted on validation set 1. Please note that the validation set comprises individuals not encountered during training, allowing for an assessment of the algorithm’s generalization capabilities. The whole data collecting process was approved and supervised by the Tsinghua University Medical Ethics Committee (Approval No.THU-01-2025-1001). A signed informed consent form was collected from each participant.
@@
Pre-training. The vision encoder is firstly pre-trained on training set using the I-JEPA method [13] with their default hyper-parameters. During our pre-training, the vision encoder is further jointly trained with the sequence-aware transformer. The training batch size is 512. The model is optimized using AdamW for 50 epochs, with the first 7 epochs designated for warm-up, followed by a cosine learning rate schedule starting at an initial learning rate of 0.00025. Data augmentation only involves random resized crop, with a crop size of $\times224$. The mask ratio is set to range between 0.3 and 0.5. The default sequence length $L$ is set to 6 unless otherwise specified.
Downstream fine-tuning. For fine-tuning, after loading the pre-trained parameters, the model is optimized using AdamW with a batch size of 512 and a learning rate of $1\times 10^{-4}$. A cosine learning rate schedule is employed, and the training is conducted for 5 epochs. For evaluation protocols (1)/(2), the vision encoder is jointly fine-tuned with other modules. Additionally, for evaluation protocol (2), we apply a drop path rate of 0.2 to the sequence-aware transformer and a layer-wise learning rate decay of 0.4 to the vision encoder. The baselines use the same settings. For evaluation protocol (3), we empirically find that freezing the vision encoder achieves better performance. All experiments are run on the Ubuntu system equipped with 8 NVIDIA RTX 3090 GPUs.
@@
Translation MAE $=\frac{1}{3}\sum_{i=1}^{3}|\mathbf{a}_{i}-\mathbf{a}^{{}^{\prime}}_{i}|,$ (11)
Rotation MAE $=\frac{1}{3}\sum_{i=4}^{6}|\mathbf{a}_{i}-\mathbf{a}^{{}^{\prime}}_{i}|.$ (12)
@@
Table 1: Comparison with baseline methods. We report mean absolute error (lower is better) on the ten most common standard planes. “Trans.” and “Rot.” represent the translation (mm) and rotation (°) errors. † indicates that our pre-trained vision encoder and initialized prediction heads are used. ‡ indicates that these methods use the same vision encoder as our method, initialized with US-IJEPA pre-trained weights.
@@
$\text{US-GuideNet}^{{\ddagger}}$ [11] 7.32 6.14 6.97 6.98 7.51 7.07 6.93 7.66 6.97 7.61 8.16 8.31 7.35 6.35 7.48 8.95 7.88 8.50 7.60 10.53 7.42 7.81
$\text{Decision-T}^{{\ddagger}}$ [36] 7.51 6.12 6.97 7.17 7.53 7.06 6.78 7.66 6.96 7.56 8.21 8.17 7.03 6.23 7.18 8.81 7.68 8.02 7.41 10.49 7.33 7.73
@@
$\text{US-GuideNet}^{{\ddagger}}$ [11] 6.26 5.71 5.48 6.10 6.02 6.19 5.39 6.39 5.34 6.21 7.02 7.22 5.85 5.58 6.00 8.18 6.24 7.46 6.11 9.43 5.97 6.85
$\text{Decision-T}^{{\ddagger}}$ [36] 6.12 5.54 5.39 5.71 5.87 6.23 5.18 6.40 5.36 6.11 7.08 6.86 6.00 5.62 6.01 8.22 6.39 7.47 6.09 9.89 5.95 6.81
@@
Table 2: Statistical significance testing of our model performance vs. the best baselines.
Method Bidirectional Unidirectional Single-frame
Trans. Rot. Trans. Rot. Trans. Rot.
Paired t-test p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001
Bootstrap (95% CI) [0.64, 0.65] [0.63, 0.64] [0.23, 0.24] [0.19, 0.20] [0.38, 0.43] [0.24, 0.29]
@@ 4.2 Comparison with Baselines
To more rigorously validate the performance improvement of our model, we conducted statistical significance tests comparing our model’s performance against that of the best baseline across three protocols. In Tab. 2, the paired t-test results show that the mean performance of our model is significantly lower than that of the baselines, while the bootstrap resampling further confirms a statistically significant improvement. Based on the above results, our model indeed shows a clear improvement over the baselines.
Table 3: Performance analysis based on demographic under bidirectional protocol.
[…]
Trans. Rot. Trans. Rot. Trans. Rot. Trans. Rot. Trans. Rot. Trans. Rot.
Decision-T 5.56 6.92 7.46 8.20 5.87 8.00 6.63 8.53 5.76 7.86 7.66 7.98
Ours 5.23 6.13 6.73 7.81 5.47 7.56 6.20 7.81 5.46 7.83 6.71 7.48
Paired t-test p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001 p<<0.0001
Bootstrap (95% CI) [0.32, 0.33] [0.78, 0.80] [0.72, 0.73] [0.38, 0.40] [0.39, 0.41] [0.44, 0.45] [0.42, 0.43] [0.71, 0.73] [0.29, 0.30] [0.02, 0.04] [0.94, 0.95] [0.49, 0.51]
4.3 Analysis on Demographic Variables
Cardiac structure is influenced by multiple demographic variables, such as disease status, sex, and age. These factors may further affect ultrasound imaging appearances and thereby introduce additional challenges to model perception and decision-making. Therefore, we conducted validation analyses on these potential influencing factors.
First, regarding disease conditions, validation set 1 includes nine subjects with cardiac diseases including atrial septal defect, ventricular septal defect, degenerative valve changes, and ventricular septal hypertrophy. The oldest subject in this group is 69 years old, with a mean age of 47.5. As shown in Tab. 1 and Tab. 2, our model outperformed the baseline model on validation set 1 containing patients. Furthermore, we tested the nine patient cases as an independent validation set. The results presented in the ”cardiac disease” column of Tab. 3 demonstrate that our model maintains stable performance while still surpassing the baseline model.
Second, we compared the performance of our proposed model with the best baseline, Decision-T, across sex and age groups on validation set 2. This validation set was specifically constructed to balance sex and age distributions, consisting of 60 subjects with an age distribution of 54.2 ± 10.7 years (max: 75, min: 26). The subjects were divided into three age groups: under 50 years, 50 to 60 years, and over than 60 years, with 20 subjects per group and a 1:1 male-to-female ratio in each. This grouping strategy ensures gender balance and a broader age coverage, thereby better reflecting the demographic characteristics of real clinical populations. In terms of sex, our model outperformed the baseline in all cases. Both models performed better in the female group, which may be attributed to the fact that approximately one-third of the male subjects were smokers—smoking can lead to lung deposits that impair imaging clarity and pose additional challenges for model perception. Overall, our model demonstrated effectiveness across both sexes and surpassed the current best model. Across age groups, our model outperformed or matched the baseline. Although some performance fluctuations were observed across age groups, these occurred in both our model and the baseline model with similar trends, likely due to inherent variations in data distribution within each subgroup. Nevertheless, our model consistently demonstrated effectiveness across all age groups and overall outperformed the current best model.
Table 4: Ablation study on sequence-aware pre-training.
Pre-train     Protocol Trans.($m$) $\downarrow$ Rot.(°) $\downarrow$     Protocol Trans.($m$) $\downarrow$ Rot.(°) $\downarrow$
✗     Unidirectional 7.41 7.90     Bidirectional 5.53 6.69
✓     7.01 (-5.4%) 7.55 (-4.4%)     5.12 (-7.4%) 6.34 (-5.2%)
@@ 4.4 Ablation and Discussion of Key Components
Protocol Trans.($m$) $\downarrow$ Rot.(°) $\downarrow$
@@
Figure 7: Ablation study under various training scanning demonstrations and sequence length.
Figure 8: Ablation of pre-training (left) and sequential modeling (right) in learning personalized features.
@@
Furthermore, since the heart undergoes continuous beating, which affects the content of cardiac ultrasound images, we further investigate the impact of the cardiac cycle on the model. Specifically, we sample 500 pairs of images from the same view but at different cardiac cycle phases, generate guidance actions toward ten standard views, and compute the average absolute error/difference between outputs from different cardiac phases. For the sequence modeling approach, we keep the same L1L-1 historical views unchanged and only varied the current input’s cardiac phase. As shown in Fig. 9, the sequential modeling method significantly improves robustness against cardiac cycle variations, whereas the single-image-based approach is heavily influenced by the cardiac phase.
@@
Table 6: Ablation study on sampling strategies. MAE values are averages of ten planes.
Strategy     Protocol Trans.($m$) $\downarrow$ Rot.(°) $\downarrow$     Protocol Trans.($m$) $\downarrow$ Rot.(°) $\downarrow$
@@
Sampling strategy. Since the scanning process generates long historical trajectories, selecting key instances from these trajectories is crucial for the model’s performance. In Tab. 6, we compare segmental sampling with two common sampling strategies: dense sampling [38] and random sampling. Dense sampling selects a sequence of L1L-1 consecutive frames preceding the current image, while random sampling selects L1L-1 frames randomly from the sampling interval to form a sequence. From the table, it can be observed that under both protocols, segmental sampling achieves the best performance, while dense sampling performs the worst, and random sampling is slightly better than dense sampling. Dense sampling performs poorly because the sampled interval is too localized, and the variation in probe poses within such a local interval is typically small. This means that the sequence contains highly localized cardiac structural information around the current plane, failing to provide global structural insights which is useful for navigating to other planes. Meanwhile, the sequences obtained through random sampling can cover a larger sampling interval, resulting in a greater variety of probe poses (spanning multiple different planes), which also means more diverse structural information. However, due to its randomness, random sampling may also face the same issue as dense sampling, where the samples could be too localized. Unlike the previous two methods, segmental sampling first divides the sampling interval into L1L-1 equal segments and then randomly samples one instance from each segment. This approach prevents samples from being confined to localized intervals, which in turn leads to an improvement in overall performance.
@@ 4.5 Ablation of Mask Hyper-parameters
Figure 11: Ablations of masking type and ratio.
@@
Figure 12: Visualization of the model’s prediction. We performed nearest-neighbor sampling from the scans to retrieve the plane most close to the model’s predictions. The Structural Similarity Index Measure (SSIM) is used to measure the content similarity between the retrieved plane and the target plane.
Figure 13: Visualization of the model’s prediction on subjects with atrial septal defect. We performed nearest-neighbor sampling from the scans to retrieve the plane most close to the model’s predictions.
@@ 4.6 Visualization
To intuitively demonstrate the model’s output, we calculate the final plane pose based on the current plane’s pose and predicted probe movement action. Using this pose, we perform nearest-neighbor retrieval from the same scan and present retrieved planes. We compare four models in Fig. 12, showing that our model’s preditions align closest with the target plane, confirming the superior precision of our guidance signal. To further validate this, we compare the structural similarity (SSIM) between the retrieved planes and the target plane, demonstrating that our model achieves the highest similarity.
To demonstrate the model’s performance on cases with cardiac conditions, we visualized the inference results for two patients with atrial septal defect in Fig. 13. The results show a high agreement between the model’s output and the expert’s annotations, confirming its potential in clinical scenarios involving cardiac diseases.
@@ 4.7 Efficiency
In ultrasound scanning, highly skilled sonographers make immediate adjustments based on the ultrasound images, a process that involves real-time decision-making. Therefore, an ultrasound guidance model must also possess real-time decision-making capabilities to provide timely guidance. To this end, we tested the decision-making time for different sequence lengths on a single RTX 3090 and A100 GPU, as shown in Tab. 7. From the table, it can be seen that our system achieves real-time inference performance on commonly used GPUs.
Table 7: Efficiency of the guidance model for a single inference.
Sequence Length 4 6 8
RTX3090 8.71 ms 10.56 ms 12.89 ms
A100 3.45 ms 4.94 ms 6.41 ms
@@ 5 Discussion and Conclusion
In this paper, we propose a cardiac probe movement guidance model which has the potential of driving autonomous robotic echocardiography or helping novice in scanning ten important standard planes. Specifically, we introduce a novel sequence-aware pre-training method to enhance the model’s ability to capture personalized cardiac structures from past sequences, addressing individual variability. Extensive experiments on a large-scale expert scanning dataset with 1.67M samples demonstrate that our model outperforms other advanced baselines and highlights the superiority of the proposed sequence-aware pre-training method.
Although experiments have validated the effectiveness of our idea, it should be noted that certain limitations remain and warrant future investigation. Firstly, the sampling strategy is manually designed, which may be a suboptimal approach. As evidenced by the ablation study, as the sequence length increases, the rate of error reduction slows down, which may be due to the presence of information redundancy or even noise in the sampled instances. Therefore, a valuable future direction is to develop methods for sampling the most informative sequences, maximizing information gain from extensive past scanning sequence data. Secondly, the current work follows an imitation learning strategy, which may be constrained by the operator’s expertise. In other words, the learned policy cannot surpass the performance of the sonographer. Given the recent advances in reinforcement learning, it is promising to explore whether such methods could learn an optimal scanning strategy. Last, information from other modalities such as ultrasound reports and sonographers’ gaze information has not yet been collected. This type of data has the potential to further enhance the model’s performance. We believe these limitations are also promising research directions that are worth further investigation.
Our work demonstrates the potential of AI-assisted ultrasound scanning, yet several challenges remain for real-world deployment. Firstly, integrating such a system into a point-of-care ultrasound platform requires substantially more diverse clinical data to achieve truly robust and clinically viable performance. Second, beyond assisting human operators, our guidance model can also drive a robotic system. However, developing such a system involves addressing critical human-robot interaction challenges—particularly how to ensure both precise execution of movements and the subject’s comfort and safety. In conclusion, while this study represents a significant step toward intelligent ultrasound assistance, translating technological innovation into real-world clinical practice demands continued interdisciplinary collaboration—integrating clinical data, AI algorithms, and even robotic control.